In a normal distribution 10% of the items are under 50 and 90% of the items are the under 80.find the values mean and standard deviation of the distribution?


1
Expert's answer
2021-06-17T12:50:32-0400

Given


"P(Z<z_1)=0.1, z_1= \dfrac{x_1-\mu}{\sigma}=\dfrac{50-\mu}{\sigma}"

"P(Z<z_2)=0.9, z_2= \dfrac{x_1-\mu}{\sigma}=\dfrac{80-\mu}{\sigma}"

"z_1\approx-1.28155, z_2\approx1.28155"


"50-\mu=-1.28155\sigma"

"80-\mu=1.28155\sigma"

"130-2\mu=0=>\mu=65"

"\sigma=\dfrac{80-65}{1.28155}\approx11.7046"